# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v1

@@ VIII-A5 Judgment and Accountability Mechanism for Agent
It is still difficult to locate and assign responsibility for the behavior of agents. For example, in a failed task execution process, it is hard to identify which steps caused the final deviation of the result, no matter they are malicious or unintentional. This is because a tiny deviation in the middle process may lead to a final gap between benign and dangerous results. Besides, it also needs a principle to quantify the responsibility for each agent or action. We believe this aspect will significant address the urgent need of the current AI ecosystem.
VIII-A6 Trade-offs between Efficiency and Accuracy
@@
High-token Communication: A larger number of tokens allows agents to convey richer contextual semantics, more detailed instructions, and more complex logic, thereby reducing ambiguity and enhancing the accuracy of multi-agent coordination. In tasks that require fine-grained understanding, verbose natural language descriptions help align goals among agents and reduce deviations. However Excessive tokens significantly increase costs and processing time, resulting in lower system efficiency and higher latency. Moreover, longer contexts expose larger attack surfaces for prompt injection and data poisoning, enabling adversaries to hide malicious content more covertly. Additionally, information overload may distract agents, causing them to infer incorrect information from irrelevant context and increasing the likelihood of hallucinations.
@@
VIII-A7 Towards Self-Organizing Agentic Networks
@@
With increasing scale of IoA, in the future, agent communication is expected to evolve toward self-organizing agentic networks, where agents autonomously discover each other, assess capabilities, negotiate collaborations, form dynamic task groups, and disband upon completion. This paradigm offers high scalability and robustness, making it well-suited for dynamic and unpredictable environments.
VIII-B Law and Regulation Aspect
@@
Apart from technical aspect, we find that there are still serious deficiencies in the laws and regulations related to agents. Such blanks cannot be cannot be remedied by techniques. We call for accelerating the improvement of laws and regulations in the following aspects.
VIII-B1 Clarify the Responsible Subject
@@
When a sold agent causes property damage or personal injury to others, it is difficult to determine the ultimate responsible subject. For example, if an intelligent robot damages the property during the execution of a task, the law-level quantification of the responsibility of the developers, user, or enterprises lacks clear definition. In addition, for problems arising from the collaborative work of multiple agents, such as an accident occurring when multiple autonomous driving vehicles are traveling in formation, there is a lack of legal provisions regarding the division of responsibilities among the enterprises to which the vehicles belong or the relevant subjects.
VIII-B2 Protect Intellectual Property Rights
@@
Nowadays, there have been a large amount of LLMs that have been open-sourced. These can act as the brain of different agents. However, even for open-source LLMs, the publishers still restrict their application scope, e.g., other developers should also open source their agents built on these LLMs. However, there still lacks laws to effectively protect such intellectual property. For example, the criteria for determining plagiarism in agents is not clear. Even if plagiarism is determined, there is still a lack of defining standards for the degree of plagiarism (e.g., 50% or 90%?). We think there urgently need related laws and regulations.
VIII-B3 Cross-border Supervision
@@
Agent communication has a transnational nature. An agent trained in one country may be used for illegal activities by people from other countries. At this time, it is difficult to determine which country’s laws apply, and there is a lack of unified international supervision standards and judicial cooperation mechanisms, which may easily lead to the difficulty of cross-border security.
@@
To our knowledge, the related formulation of laws and regulations (such as those related to agent crimes) lag far behind the development of agents. For example, how to define the theft and misappropriation of agents, the accident responsibility of autonomous driving agents
IX Conclusion
@@
This survey systematically reviews the security issues of agent communication. We first highlight the differences between previous related surveys and this survey, and summarize the preliminaries of LLM-driven agents. Then, we make a definition and classification of agent communication to help future researchers to quickly classify and evaluate their work. Next, we detailed illustrate the communication protocols, security risks, and possible defense countermeasures for three agent communication stages, respectively. Finally, we discuss the open issues and future directions from technical and legal aspects, respectively.

### v2

@@ IX-A5 Judgment and Accountability Mechanism for Agent
It is still difficult to locate and assign responsibility for the behavior of agents. For example, in a failed task execution process, it is hard to identify which steps caused the final deviation of the result, no matter they are malicious or unintentional. This is because a tiny deviation in the middle process may lead to a final gap between benign and dangerous results. Besides, it also needs a principle to quantify the responsibility for each agent or action. We believe this aspect will significantly address the urgent need of the current AI ecosystem.
IX-A6 Trade-offs between Efficiency and Accuracy
@@
High-token Communication: A larger number of tokens allows agents to convey richer contextual semantics, more detailed instructions, and more complex logic, thereby reducing ambiguity and enhancing the accuracy of multi-agent coordination. In tasks that require fine-grained understanding, verbose natural language descriptions help align goals among agents and reduce deviations. However, excessive tokens significantly increase costs and processing time, resulting in lower system efficiency and higher latency. Moreover, longer contexts expose larger attack surfaces for prompt injection and data poisoning, enabling adversaries to hide malicious content more covertly. Additionally, information overload may distract agents, causing them to infer incorrect information from irrelevant context and increasing the likelihood of hallucinations.
@@
IX-A7 Towards Self-Organizing Agentic Networks
@@
With the increasing scale of IoA, in the future, agent communication is expected to evolve toward self-organizing agentic networks, where agents autonomously discover each other, assess capabilities, negotiate collaborations, form dynamic task groups, and disband upon completion. This paradigm offers high scalability and robustness, making it well-suited for dynamic and unpredictable environments.
IX-B Law and Regulation Aspect
@@
Apart from the technical aspect, we find that there are still serious deficiencies in the laws and regulations related to agents. Such blanks cannot be remedied by techniques. We call for accelerating the improvement of laws and regulations in the following aspects.
IX-B1 Clarify the Responsible Subject
@@
When a sold agent causes property damage or personal injury to others, it is difficult to determine the ultimate responsible subject. For example, if an intelligent robot damages the property during the execution of a task, the law-level quantification of the responsibility of the developers, users, or enterprises lacks a clear definition. In addition, for problems arising from the collaborative work of multiple agents, such as an accident occurring when multiple autonomous driving vehicles are traveling in formation, there is a lack of legal provisions regarding the division of responsibilities among the enterprises to which the vehicles belong or the relevant subjects.
IX-B2 Protect Intellectual Property Rights
@@
Nowadays, there has been a large number of LLMs that have been open-sourced. These can act as the brain of different agents. However, even for open-source LLMs, the publishers still restrict their application scope, e.g., other developers should also open-source their agents built on these LLMs. However, there still lacks laws to effectively protect such intellectual property. For example, the criteria for determining plagiarism in agents is not clear. Even if plagiarism is determined, there is still a lack of defining standards for the degree of plagiarism (e.g., 50% or 90%?). We think there urgently need related laws and regulations.
IX-B3 Cross-border Supervision
@@
Agent communication has a transnational nature. An agent trained in one country may be used for illegal activities by people from other countries. At this time, it is difficult to determine which country’s laws apply, and there is a lack of unified international supervision standards and judicial cooperation mechanisms, which may easily lead to difficulties in cross-border security.
@@
To our knowledge, the related formulation of laws and regulations (such as those related to agent crimes) lags far behind the development of agents. For example, how to define the theft and misappropriation of agents, the accident responsibility of autonomous driving agents
TABLE VIII: Abbreviation Table.
Abbr. Full Form
A2A Ageng-to-Agent Protocol
ACN Agent Communication Network
ACP-AGNTCY Agent Connect Protocol by AGNTCY
ACP-IBM Agent Communication Protocol by IBM
ACP-AgentUnion Agent Communication Protocol by AgentUnion
AI artificial intelligence
AITP Agent Interaction & Transaction Protocol
ANP Agent Network Protocol
API Application Programming Interface
CNN Convolutional Neural Network
CoT Chain of Thought
DID Decentralized Identifier
DNS Domain Name System
DoS Denial of Service
FNN Feedforward Neural Network
GAN Generative Adversarial Network
GNN Graph Neural Network
IoA Internet of Agents
LMOS Language Model Operating System Protocol
LLM Large Language Model
LOKA Layered Orchestration for Knowledgeful Agents
LSTM Long Short-Term Memory
MAS Multi-Agent Systems
MCP Model Context Protocol
MITM Man-in-the-Middle
OAuth2 Open Authorization 2.0
PXP PXP protocol
RAG Retrieval-Augmented Generation
RNN Recurrent Neural Network
SSL Secure Sockets Layer
SPPs Spatial Population Protocols
TLS Transport Layer Security
X Conclusion
@@
This survey systematically reviews the security issues of agent communication. We first highlight the differences between previous related surveys and this survey, and summarize the preliminaries of LLM-driven agents. Then, we make a definition and classification of agent communication to help future researchers quickly classify and evaluate their work. Next, we detailedly illustrate the communication protocols, security risks, and possible defense countermeasures for three agent communication stages, respectively. Then, we conduct experiments using MCP and A2A to help illustrate the new attack surfaces brought by agent communication. Finally, we discuss the open issues and future directions from technical and legal aspects, respectively.